A 10.8 kg object is moving along the + x axis with a speed of 10.0 m/s . A constant force is applied to this object which gradually brings the object to rest . The object traveled 13.0 meters during this time . To the nearest 0.1 N , what was the magnitude of the force ? Do not include units in your answer .


1
Expert's answer
2021-02-19T10:31:06-0500

The acceleration provided by the force is:


"a= \dfrac{v_f-v_i}{t} = -\dfrac{v_i}{t}"

where "v_f = 0m\/s" is the final speed of the object, "v_i = 10m\/s" is the initial speed, "t" is the time the object comes to rest.

The distance travelled in this time is given as follows:


"d = v_i t + \dfrac{at^2}{2}\\
d = v_i t - \dfrac{v_it}{2} = \dfrac{v_it}{2}"

Since "d = 13m", the time will be:


"t = \dfrac{2d}{v_i}"

Then the magnitude of the acceleration will be:


"a = \dfrac{v_i}{t} = \dfrac{v_i^2}{2d}"

Then the force will be (according to the second Newton's law):


"F = ma"

where "m = 10.8kg" is the mass of the object. Thus, obtain:


"F = \dfrac{mv_i^2}{2d}\\
F = \dfrac{10.8\cdot 10^2}{2\cdot 13}\approx 41.5N"

Answer. 41.5 N.